OpenStudy (anonymous):

2x-5(x-3)=2(x-10) please show work.

OpenStudy (saifoo.khan):

2x-5(x-3)=2(x-10) 2x - 5x + 15 = 2x - 20

OpenStudy (saifoo.khan):

Can you solve it now

OpenStudy (anonymous):

no, this is the part that i got confused on.

OpenStudy (saifoo.khan):

Ok, let's remove your confusion! xD 2x - 5x + 15 = 2x - 20 Add like terms, (2-5)x + 15 = 2x - 20 -3x + 15 = 2x - 20 Do you get till here?

OpenStudy (anonymous):

yees.

OpenStudy (saifoo.khan):

Cool, now let's shuffle. -3x + 15 = 2x - 20 we brought 2x on left side, so it will change it's sign. Same as the case with 15. -3x - 2x = -20 - 15

OpenStudy (saifoo.khan):

Do you get this part?

OpenStudy (anonymous):

yeep.

OpenStudy (saifoo.khan):

Let's proceed, WAIT. this time, you try it.. otherwise i willl help! ^_^

OpenStudy (anonymous):

im not really sure.

OpenStudy (saifoo.khan):

try it, no matter if it's incorrect.

OpenStudy (anonymous):

would you add 20 to each side and then 15?

OpenStudy (anonymous):

or would you do -3-2 (which are x) and then -20-15?

OpenStudy (saifoo.khan):

No. now as we have to solve it.. we have to add like terms always. so, the numbers with variables with numbers with variables; -3x & -2x & numbers without variables with numbers without variables. like: -20 & -15

OpenStudy (saifoo.khan):

Yes, you are doing it the right way,,, try it..

OpenStudy (anonymous):